OpenStudy (anonymous):

What is the slope of the line whose equation is 3x+y=4? a)3x b) -3 c) 3 d) 4

OpenStudy (anonymous):

try the formula

OpenStudy (anonymous):

im pretty sure its C but im not that great with this stuff..

OpenStudy (anonymous):

3 its obviously 3 using y=mx+b where m is the slope in the equation 3 is the slope

OpenStudy (anonymous):

both are incorrect. try bringing the equation to the slope intercept form. y = mx + b

OpenStudy (anonymous):

subtract 3x from both sides, giving y = -3x+4

OpenStudy (anonymous):

ohhh sorry its -3 y=4-3x